OpenStudy (anonymous):

I need serious help. MEDALS AWARDED AND FAN ADD. I need help with figuring out the step-by-step process of writing f(x) = x^2 - 5x + 6 in general vertex form. Please help!

OpenStudy (anonymous):

@campbell_st @ranga @Calliope

OpenStudy (math&ing001):

f(x)= x^2 - 5x + 6 = x^2 - 2 * 5/2 * x + (5/2)^2 - (5/2)^2 + 6 = (x-5/2)^2 - 25/4 + 6 = (x-5/2)^2 -1/4

OpenStudy (anonymous):

@math&ing001 could you please explain what you did? I need help with step by step stuff. Thank you!

OpenStudy (btaylor):

First, you complete the square on the x^2 + 5x part to give you x^2 + 2(5/2)x + 25/4. Then, since you added 25/4, you have to subtract 25/4 to keep the same equation. So, you have (x+5/2)^2 -25/4 + 6. This gives you (x-5/2)^2 - 1/4. Do you understand?

OpenStudy (anonymous):

Yes, thank you soooo much! @BTaylor I'll message you if I need more help, if that is okay with you(:

OpenStudy (btaylor):

Sure!

OpenStudy (anonymous):

and thank you @math&ing001

OpenStudy (math&ing001):

You're welcome !